OpenStudy (anonymous):

Help plz Choose the equation that could be used to find three consecutive integers whose sum is 36. n + (n + 2) + (n + 4) = 36 n + (n + 1) + (n + 3) = 36 n + (n + 1) + (n + 2) = 36 n + (n − 1) + (n − 3) = 36

OpenStudy (anonymous):

Choose the equation that could be used to find three consecutive integers whose sum is 36. n + (n + 2) + (n + 4) = 36 n + (n + 1) + (n + 3) = 36 n + (n + 1) + (n + 2) = 36 n + (n − 1) + (n − 3) = 36